OpenStudy (anonymous):

if a =2 b=3 and c=5find the value of the folowing 4a+4b-c whats that?

OpenStudy (anonymous):

4(2)+4(3)-5 =8+12-5 =15

OpenStudy (anonymous):

done by substitution, sub in the values you have for a,b and c into the equation. then solve:)

OpenStudy (anonymous):

thank you